# Mirror Descent Policy Optimisation for Robust Constrained Markov Decision Processes

### v1

@@ 4.2 Global optimality of Lagrangian TMA
4.3 Global optimality of the Lagrangian policy update in RCMDPs
Since at each loop of TMA, it is possible to obtain arbitrary precision, we now turn to proving the robust objective regret upper bound for the policy $\pi_{\theta}$ in the softmax parametrisation. Our technique for RCMDPs is similar to the results for RMDPs (Wang and Petrik,, 2024), as our proof is based on finding a correspondence between the robust objective and the nominal objective and then making use of traditional MDP results by Mei et al., (2020).
4.3.1 Lagrangian policy gradient
To extend the analysis of RMDPs and traditional MDPs to the constrained setting, we first develop a few key lemmas for Lagrangian-based policy gradient. These results will then be used for analysing the convergence of RCMDPs under squared error and KL-divergence based Bregman divergences.
Lemma 3 (Policy gradient for softmax parametrisation ).
Let $\pi$ be a policy parametrised according to
$\pi_{\theta}(a|s)=\frac{\exp(\theta_{s,a})}{\sum_{a^{\prime}\in\mathcal{A}}\theta_{s,a^{\prime}}}$
and let $\mathbf{A}_{\pi_{\theta}}(s,a)=\mathbf{Q}_{\pi_{\theta}}(s,a)-\mathbf{V}_{\pi_{\theta}}(s)$ be the advantage function. Then the Lagrangian policy gradient is given by
$\frac{\partial\mathbf{V}_{\pi_{\theta}}(\rho)}{\partial\theta_{s,a}}=\frac{1}{1-\gamma}d_{\rho}^{\pi,p}(s)\pi_{\theta}(a|s)A_{\pi_{\theta}}(s,a)\,.$ (21)
@@ 4.3.1 Lagrangian policy gradient
The proof follows from Eq. 10 in Agarwal et al., (2021) and by considering the Lagrangian as a value function. ∎
4.3.2 Regret of mirror descent policy optimisation with robust Lagrangian
When using the squared error as a distance-generating function, the mirror descent update is equivalent to traditional gradient descent (see Table 5). This can be exploited to derive results for RCMDPs. In particular, using smoothness and continuity properties, an equivalent gradient descent sequence can be derived to provide guarantees based on traditional gradient descent over a traditional MDP.
We first restate the supporting lemma from Wang and Petrik, (2024).
Lemma 4 (Existence of equivalent gradient descent sequence, Lemma B.5 in Wang and Petrik, (2024)).
For any $t^{\prime}\geq 0$, any $\theta\in\mathbb{R}^{d}$, and any optimisation problem $\min_{\theta\in\mathbb{R}}f(\theta)$ in which $f:\mathbb{R}^{d}\to\mathbb{R}$ is $L_{\theta}$-Lipschitz and $l_{\theta}$-smooth, there exists a sequence $\left\{\theta_{t}\right\}_{t\geq 0}$, generated by the gradient descent update with constant learning rate $\eta>0$
$\theta_{t+1}=\theta_{t}-\eta\nabla f(\theta_{t})\,,$ (22)
such that $\theta_{t^{\prime}}=\theta$.
Without entropy regularisation
Using Lemma 4, we derive an RCMDP analogue to Theorem 5.5 in Wang and Petrik, (2024). It can be applied directly to softmax parametrisations with squared error Bregman divergence and indirectly to the KL-divergence Bregman divergence as shown later in Section 4.4. We first state three important properties that are essential to the proof of the convergence rate.
First, the Lagrangian function has essential smoothness and continuity properties.
Lemma 5 (Smooth and continuous Lagrangian).
Under the softmax parametrisation, the Lagrangian $\mathbf{V}(\theta)$ with multipliers $\lambda\in\mathbb{R}^{m}$ is $l_{\theta}$-smooth and $L_{\theta}$-Lipschitz continuous with
$l_{\theta}(\lambda)=\frac{16F_{\lambda}}{(1-\gamma)^{3}}$ (23)
@@ 4.3.2 Regret of mirror descent policy optimisation with robust Lagrangian
$L_{\theta}(\lambda)=\frac{2\sqrt{2}F_{\lambda}}{(1-\gamma)^{2}}\,.$ (24)
@@
From Lemma 4.4 in Wang and Petrik, (2024), the value function based on costs in $[0,1]$ has $L_{\theta}=\frac{\sqrt{2}}{(1-\gamma)^{2}}$ and $l_{\theta}=\frac{8}{(1-\gamma)^{3}}$. Since the Lagrangian is based on costs and constraint-costs in $[-1,1]$, the result follows after using the normalisation by $2F_{\lambda}$ (see Eq. 11 and 12. ∎
Second, we note that under a fixed set of Lagrangian multipliers, standard convergence rate results for softmax policies can be used.
Lemma 6 (Convergence rate for softmax policies for Lagrangian gradient descent).
For Lagrangian gradient descent over logits, and Lemma 5, with settings of $M=\inf_{p\in\mathcal{P},\pi\in\Pi}M_{p}(\pi)$, $U=\inf_{p\in\mathcal{P}}U_{p}$, fixed $\lambda_{k}=\lambda_{k-1}=\lambda$ for all $j\in[m]$, and $\eta=\frac{1}{l_{\theta}((\lambda)}$ according to Theorem 5 that for a fixed transition kernel $p\in\mathcal{P}$,
$\mathbf{V}_{\pi_{\theta_{t}},p}(\rho)-\mathbf{V}_{\pi_{\theta^{*}},p}(\rho)\leq\left\lVert\frac{1}{\mu}\right\rVert_{\infty}\frac{32SM^{2}F_{\lambda}}{U_{p}(1-\gamma)^{6}t}\,.$ (25)
As a third and last auxiliary result, we note that the RCMDP objective in Eq. 3 can be expressed in terms of the optimal (adversarial) transition dynamics at macro-iteration $k$ can be expressed in terms of the transition dynamics found by LTMA and its error tolerance.
Lemma 7.
Let $\Phi$ be the objective in Eq. 3 and $\pi^{*}$ its optimal solution. Then for any macro-step $k\geq 0$,
$\Phi(\pi_{k})-\Phi(\pi^{*})$ $\leq\mathbf{V}_{\pi_{k},p_{k}}(\rho)+\epsilon_{k}^{\prime}-\mathbf{V}_{\pi^{*},p_{k}}(\rho)\,,$ (26)
where $\epsilon_{k}^{\prime}>0$ is the error tolerance for LTMA at iteration $k$.

### v2

@@ 5.2 Global optimality of Lagrangian TMA
5.3 Analysis of Approximate TMA
While the above-mentioned LTMA algorithm was formulated in terms of exact policy gradient evaluations, such computations are intractable in practice. Now we turn to the design of an algorithm for TMA based on approximate policy gradient evaluations, which we call Approximate TMA (see Algorithm 1). Since it has not been proposed in earlier work, the algorithm will be presented in a generic manner, i.e. for general value functions. To maintain consistency across different value functions (including the Lagrangian and traditional value functions), we deal with their difference through a factor $C>0$ that scales the maximal cost at each time step.
To construct the Approximate TMA algorithm, the definition of the action next-state value function provides a useful short hand that is an essential part of the transition gradient for a directly parametrised transition kernel.
Definition 2 (Action next-state value function (Definition 3.2 in Li et al., (2023))).
For any cost function $c:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ and starting distribution $\rho$, the action next-state value function for transition kernel $p$ and policy $\pi$ is defined as
$G_{\pi,p}(s,a,s^{\prime})=\mathbb{E}\left[\sum_{l=0}^{\infty}\gamma^{l}c(s_{l},a_{l},s_{l+1})|s_{0}=s,a_{0}=a,s_{1}=s^{\prime},\pi,p\right]\,.$ (24)
It can also be written as $G_{\pi,p}(s,a,s^{\prime})=c(s,a,s^{\prime})+\gamma V_{\pi,p}(s^{\prime})$ (Wang et al.,, 2024).
The directly parametrised transition kernel can now be written in terms of Definition 2 based on the lemma below.
Lemma 3 (Partial derivative for approximate TMA).
Under the direct parametrisation, the partial derivative of $\hat{V}_{\pi,p}(\rho)$ has the form
$\frac{\partial}{\partial_{p_{s,a,s^{\prime}}}}\hat{V}_{\pi,p}(\rho)=\frac{1}{1-\gamma}d_{\rho}^{\pi,p}(s)\pi(a|s)\hat{G}_{\pi,p}(s,a,s^{\prime})\,.$ (25)
@@ 5.3 Analysis of Approximate TMA
The proof follows directly by applying Lemma 5.1 of Wang et al., (2024) to the approximate value function. ∎
The resulting update rule is then given by
$p^{t+1}(\cdot|s,a)=\mathrm{arg}\max_{p\in\Delta(\mathcal{S})}\eta_{p}(t)\frac{1}{1-\gamma}d_{\rho}^{\pi,p^{t}}(s,a)\hat{G}_{\pi,p^{t}}(s,a,\cdot)-\alpha_{p}B_{d_{\rho}^{\pi,p^{t}}}(p,p^{t})\quad\text{for all }\mathcal{S}\times\mathcal{A}\,.$ (26)
As a consequence of this update rule, the value difference in $p^{t+1}$ and $p^{t}$ can be bounded. In particular, the ascent property of TMA (Lemma 5.4 of Wang et al., (2024)) is a key result that can be used to construct the proof for the approximate TMA algorithm. The property, which is given below, notes that conditioning on the state or state-action, a pushback property can be achieved, and this can be used to show that (under perfect gradient information) TMA updates always result in increasing value.
Lemma 4 (Ascent property of TMA (Lemma 5.4 of Wang et al., (2024))).
Denoting $G_{\pi,p}(s,\cdot,\cdot)=\left(\pi(a|s)G_{\pi,p}(s,a,\cdot)\right)_{a\in\mathcal{A}}$, for any $p\in\mathcal{P}_{s}$ and $\pi\in\Pi$, we have that
$\langle G_{\pi,p}(s,\cdot,\cdot),p-p^{t+1}(\cdot|s,\cdot)\rangle+B(p^{t+1}(\cdot|s,\cdot),p^{t}(\cdot|s,\cdot))\leq B(p,p^{t}(\cdot|s,\cdot))-B(p,p^{t+1}(\cdot|s,\cdot))\,$, (27)
where $B(p,q)$ is the unweighted Bregman divergence between $p$ and $q$. Moreover,
$V_{\pi,p^{t}}(\rho)-V_{\pi,p^{t+1}}(\rho)\leq 0\,.$ (28)
With imperfect information on the gradient, a monotonic increase in value cannot be guaranteed. However, it is possible to bound the potential decrease, resulting in the following ascent property of Approximate TMA.
Lemma 5 (Ascent property of Approximate TMA).
Let $t\geq 0$ be the current iteration, let $V:\mathcal{S}\to\mathbb{R}$ be a value function with instantaneous cost bounded in [C,C][-C,C], and let $\pi\in\Pi$. Suppose we have an estimate $\hat{G}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ of the action next-state value function such that
$\left\lVert G_{\pi,p^{t}}-\hat{G}_{\pi,p^{t}}\right\rVert_{\infty}\leq\epsilon^{\prime}\,.$ (29)
Then at iteration $t+1$, Approximate TMA satisfies
$V_{\pi,p^{t}}(\rho)-V_{\pi,p^{t+1}}(\rho)\leq\frac{2C\epsilon^{\prime}}{1-\gamma}\,.$ (30)
Proof.
Using the first performance difference lemma (Lemma 22) and the ascent property (Eq. 28), we have that
$V_{\pi,p^{t}}(\rho)-V_{\pi,p^{t+1}}(\rho)$ $=\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)\rangle$ (Lemma 21)
$=\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\left(\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle+\right$.
$\hskip 113.81102pt\left.\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)-\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle\right)$ (decomposing)
$\leq\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\langle p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a),G_{\pi,p^{t}}(s,a,\cdot)-\hat{G}_{\pi,p^{t}}(s,a,\cdot)\rangle$ (Eq. 28)
$\leq\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\left\lVert G_{\pi,p^{t}}(s,a,\cdot)-\hat{G}_{\pi,p^{t}}(s,a,\cdot)\right\rVert_{\infty}\left\lVert p^{t}(\cdot|s,a)-p^{t+1}(\cdot|s,a)\right\rVert_{1}$ (Hölder’s inequality)
$\leq\frac{1}{1-\gamma}\sum_{s\in\mathcal{S}}d_{\rho}^{\pi,p^{t+1}}(s)\sum_{a\in\mathcal{A}}\pi(a|s)\left\lVert G_{\pi,p^{t}}(s,a,\cdot)-\hat{G}_{\pi,p^{t}}(s,a,\cdot)\right\rVert_{\infty}\left(\left\lVert p^{t}(\cdot|s,a)\right\rVert_{1}+\left\lVert p^{t+1}(\cdot|s,a)\right\rVert_{1}\right)$ (triangle inequality)
$\leq\frac{2C\epsilon^{\prime}}{1-\gamma}\,.$ (Eq. 29)
With the ascent property established, the theorem below bounds the regret of Approximate TMA. In the context of Robust Sample-based PMD-PD, the following algorithm can be used to obtain the updated transition kernel $p_{k+1}$ such that it closely matches $p_{k}^{*}$, the worst-case transition kernel for a particular macro-iteration, using the Lagrangian as a value function. However, since it is also of independent interest, it is formulated more generally here with respect to any policy and independent of the macro-iteration.
Theorem 1 (Regret bound for Approximate TMA).
Let $V:\mathcal{S}\to\mathbb{R}$ be a value function with instantaneous cost bounded in [C,C][-C,C], fix $\pi\in\Pi$, and let $p^{*}\in\mathrm{arg}\min_{p\in\mathcal{P}}V_{\pi,p}(\rho)$. Suppose we have an estimate $\hat{G}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\to\mathbb{R}$ of the action next-state value function which is $\epsilon^{\prime}$-precise as in Eq. 29, and the step sizes follow $\eta_{p}(0)\geq\alpha_{p}(0)\frac{\gamma}{C(1-\gamma)}B_{d_{\rho}^{\pi,p^{0}}}(p^{*},p^{0})$, $\eta_{p}(t)=\eta_{p}(t-1)/\gamma$ and $\alpha_{p}(t)=\alpha_{p}(0)>0$ for all $t\geq 1$. It follows that approximate TMA (Algorithm 1) yields a regret bounded by
$V_{\pi,p^{*}}(\rho)-V_{\pi,p^{t}}(\rho)\leq\left(1-\frac{1}{M}\right)^{t}\frac{3C}{1-\gamma}+\frac{4M}{1-\gamma}\epsilon^{\prime}\,.$ (31)
for all $t\geq 1$, where $M$ is given by Eq. 18.
Proof.
The proof decomposes the approximate G-value into the true G-value and the estimation error, which allows using the ascent property of TMA and a limited factor $\mathcal{O}(\epsilon^{\prime})$, followed by further derivations based on pushback properties. The full proof is given in Appendix C.2. ∎
Algorithm 1 Approximate TMA.
1:Inputs: Discount factor $\gamma\in[0,1)$, sample size $M_{G,t}$, episode lengths $N_{G,t}$, and learning rate schedule $\eta_{p}(t)$ for all $t\in\{0,\dots,t_{k}^{\prime}-1\}$, penalty coefficient $\alpha_{p}>0$, cost function $c:\mathcal{S}\times\mathcal{A}\to\mathbb{R}$, initial transition kernel $p^{0}$, and fixed policy $\pi$.
2:for $t=0,\dots,t_{k}^{\prime}-1$ do
3:  Generate $M_{G,t}$ samples of length $N_{G,t}$ based on $\pi$ and $p^{t}$.
4:  if using function approximation then
5:   $\hat{G}_{\pi,p^{t}}\leftarrow\mathrm{arg}\min_{G\in\mathcal{F}_{G}}\frac{1}{M_{G,t}}\sum_{j=1}^{M_{G,t}}\left(G(s_{0}^{j},a_{0}^{j},s_{1}^{j})-\sum_{l=0}^{N_{G,t}-1}\gamma^{l}\left[c(s_{l}^{j},a_{l}^{j},s_{l+1}^{j})\right]\right)^{2}$.
6:  else
7:   for $(s,a,s^{\prime})\in\mathcal{S}\times\mathcal{A}\times\mathcal{S}$ do
8:     $\hat{G}_{\pi,p^{t}}(s,a,s^{\prime})=\frac{1}{M_{G,t}}\sum_{j=1}^{M_{Q,k}}\sum_{l=1}^{N_{G,t}-1}\gamma^{l}c(s_{l}^{j},a_{l}^{j},s_{l+1}^{j})$ where $s_{0}=s,a_{0}=a,s_{1}=s^{\prime}$.
9:   end for
10:  end if
11:  if using direct parametrisation then
12:   $p^{t+1}(\cdot|s,a)=\mathrm{arg}\max_{p\in\Delta(\mathcal{S})}\eta_{p}(t)\left\langle p,\hat{G}_{\pi,p^{t}}(s,a,\cdot)\right\rangle-\alpha_{p}B_{d_{\rho}^{\pi,p^{t}}}(p,p^{t})$ for all $\mathcal{S}\times\mathcal{A}$.
13:  else
14:   $\xi^{t+1}\leftarrow\mathrm{arg}\max_{\xi\in\mathcal{U}}\eta_{p}(t)\left\langle\nabla_{\xi}\hat{V}_{\pi,p^{t}}(\rho),\xi\right\rangle-\alpha_{p}B_{d_{\rho}^{\pi,p^{t}}}(\xi,\xi^{t})$. $\triangleright$ Generic form
15:   Define $p^{t+1}:=p_{\xi^{t+1}}$.
16:  end if
17:end for
Theorem 1 relies on an $\epsilon$-precise estimate of the action next-state value, which can be established within $\mathcal{O}(\epsilon^{2})$, as shown in the lemma below.
Lemma 6 (Approximate action next-state value).
Let $k\geq 0$, let $\delta\in(0,1)$ and let $t\in\{0,\dots,t_{k}^{\prime}-1\}$. Moreover, let $C>0$ be the maximal absolute value of the instantaneous cost. Then provided the number of trajectories for the estimator $\hat{G}$ (l.7–9 of Algorithm 1) satisfies
$M_{G,t}\geq\frac{2\gamma^{-2N_{G,t}}}{(1-\gamma)^{2}}\log\left(\frac{2S^{2}At_{k}^{\prime}}{\delta}\right)\,$,
Algorithm 1 provides an approximately correct estimate such that with probability at least $1-\delta$,
$\left\lVert\hat{G}_{\pi,p^{t}}-G_{\pi,p^{t}}\right\rVert_{\infty}\leq 2C\frac{\gamma^{N_{G,t}}}{1-\gamma}\quad\text{for all }t\in\{0,\dots,t_{k}^{\prime}-1\,.$ (32)
Proof.
The proof of this lemma is based on Hoeffding’s inequality and is given in Appendix C.3. ∎
Due to the above results, the sample complexity of Approximate TMA can be bounded by $\tilde{\mathcal{O}}\left(\epsilon^{-2}\right)$.
Lemma 7 (Approximate TMA sample complexity).
Let $k\geq 0$, Let $\epsilon^{\prime}>0$, $\delta\in(0,1)$, and define $M$ according to Eq. 18. Moreover, let $C>0$ be the maximal absolute value of the instantaneous cost, and let
$t_{k}^{\prime}\geq\log_{\frac{M}{M-1}}\left(\frac{6C}{\epsilon^{\prime}(1-\gamma)}\right)\,.$
Define step sizes $\eta_{p}(0)\geq\frac{\gamma}{1-\gamma}B_{d_{\rho}^{\pi,p^{*}}}(p^{*},p^{0})$, $\eta_{p}(t)=\eta_{p}(t-1)/\gamma$ and $\alpha_{p}(t)=\alpha_{p}(0)$ for all $t\in[t_{k}^{\prime}]$. Moreover, let
$N_{G,t}=H\geq\log_{\frac{1}{\gamma}}\left(\frac{16MC}{(1-\gamma)^{2}\epsilon^{\prime}}\right)$
@@
$M_{G,t}=M_{G}=\frac{2\gamma^{-2H}}{(1-\gamma)^{2}}\log\left(\frac{2S^{2}At_{k}^{\prime}}{\delta}\right)$
for all $t\in[t_{k}^{\prime}]$. Then Algorithm 1 yields a transition kernel $p^{t_{k}^{\prime}}$ such that
$V_{\pi,p^{*}}(\rho)-V_{\pi,p^{t_{k}^{\prime}}}(\rho)\leq\epsilon^{\prime}$ (33)
within at most
$n=\tilde{\mathcal{O}}\left(\frac{S^{2}AM^{3}}{(1-\gamma)^{4}\epsilon^{\prime 2}}\right)\,$ (34)
samples.
@@
The proof bounds both parts in the RHS of Eq. 31 with $\epsilon^{\prime}/2$ and then counts the number of state-action-state triples, the number of iterations, the horizon, and the number of trajectories per mini-batch. The full proof is given in Appendix C.4. ∎
5.4 Analysis of Robust Sample-based PMD-PD
Having shown results Lagrangian TMA in oracle-based and sample-based settings, we now turn to proving the sample complexity of Robust Sample-based PMD-PD (Algorithm 2). The algorithm modifies Sample-based PMD-PD (Liu et al., 2021b, ) by accounting for the mismatch between the nominal transition kernel and the optimal transition kernel in the min-max problem in Eq. 4. The complexity analysis below is slightly simplified by dropping purely problem-dependent constants $\gamma$, $\eta$, $\zeta$ from the big-O statements.
As a first auxiliary result, we note that the robust Lagrangian objective in Eq. 4 can be upper bounded based on the Lagrangian at macro-iteration $k$, obtained after the policy update and the LTMA update, and two error terms, namely the LTMA error tolerance and the suboptimality of the dual variable. These terms can then be bounded using Lemma 2, Lemma 31, and Lemma 28.
Lemma 8 (Upper bound on robust Lagrangian regret).
Let $\Phi$ be the objective in Eq. 4 and let $\pi^{*}\in\mathrm{arg}\inf_{\pi\in\Pi}\Phi(\pi)$. Then for any macro-step $k\geq 1$,
$\Phi(\pi_{k})-\Phi(\pi^{*})$ $\leq\mathbf{V}_{\pi_{k},p_{k}}(\rho;\lambda_{k-1})+\epsilon_{k}^{\prime}+\Delta_{\lambda_{k}}-\mathbf{V}_{\pi^{*},p_{k}}(\rho;\lambda_{k-1})\,$, (35)
where $\epsilon_{k}^{\prime}>0$ is the error tolerance for LTMA, and $\Delta_{\lambda_{k}}=\sum_{j=1}^{m}(\lambda_{k-1,j}^{*}-\lambda_{k-1,j})V_{\pi_{k},p_{k-1}^{*}}^{j}(\rho)$ where $(p_{k-1}^{*},\lambda_{k-1}^{*})=\mathrm{arg}\max_{p\in\mathcal{P},\lambda\geq 0}\mathbf{V}_{\pi_{k},p}(\rho;\lambda)$.